OpenStudy (anonymous):

Pls help.. Andy has two similar conical hats, Hat A and Hat B. The diameter of Hat A is 18 inches, and the diameter of Hat B is 9 cm. What is the ratio of the volume of Hat A to the volume of Hat B?

OpenStudy (johnweldon1993):



OpenStudy (johnweldon1993):

\[\large V = \frac{1}{3}\pi r^2 h\] Now we can only assume that the heights of the hats are the same...so in our ratio \[\large \frac{1}{3}\pi r^2 h : \frac{1}{3}\pi r^2 h\] the 'h' will cancel being they are the same...as a matter of fact...everything but the radius is the same... Hat A has a radius of 9 inches..and Hat B has a radius of 1.77 inches (notice how we are given the radius in CM not IN...so since those are the only differences...the ratio will simply be \[\large 9 : 1.77\]